OpenStudy (anonymous):

Determine the horizontal and vertical asymptotes for the following functions: a) g(x) = sin(x-pi/4) b) h(x) = sin x/2 c) j(x) = 2 sin x - 1

OpenStudy (anonymous):

@Hero @lgbasallote @experimentX plz help guys !

OpenStudy (dumbcow):

asymptotes for sin and cos functions ??

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

they are never undefined and functions never approach infinity....thus vertical asymptote does not exist

OpenStudy (experimentx):

for such kinda function, you have envelopes than asymptotes

OpenStudy (anonymous):

i want to know only about asymptotes

OpenStudy (anonymous):

the horizontal and vertical?

OpenStudy (dumbcow):

are you sure you don't want to know about periods and phase shifts ?

OpenStudy (anonymous):

only asymptotes ... i know period and phase shifts

OpenStudy (experimentx):

http://www.sagemath.org/calctut/inflimits.html

OpenStudy (experimentx):

info about oscillating functions is below

OpenStudy (anonymous):

ok .. what about your opinions regarding asymptotes ... the knowledge which you know regarding them?

OpenStudy (experimentx):

I don't think it would have one.

OpenStudy (anonymous):

none of those functions have vertical asymptotes..... there is no "c" such that \(\large \lim_{x \to \infty } f(x) = \pm\infty\)

OpenStudy (anonymous):

likewise for horizontal asymptotes, there is no "c" such that: \(\large \lim_{x \to \infty}f(x)=c \)

OpenStudy (experimentx):

\[\large \lim_{x \to c} f(x) = \pm\infty\]

OpenStudy (anonymous):

i could never get that "c apporoach infinity" under the limit... thanks X.... :)

OpenStudy (anonymous):

@ hero?

OpenStudy (anonymous):

@Hero ... help

OpenStudy (anonymous):

then?

OpenStudy (zzr0ck3r):

then see there is no asymptote:)

OpenStudy (anonymous):

soorry but i can't graph on calculator

OpenStudy (anonymous):

just tell me from basic definition

OpenStudy (anonymous):

did u know it or not ?

OpenStudy (anonymous):

then how can i find for sin and cosine functions?

OpenStudy (zzr0ck3r):

graph or think about 1/x note x can never equal 0 but it can get as close as it does, and as its closer and closer 1/x shoots up to infinity. ps you can graph things on the internets for free thus no need to own a calculator if you have the interwebs and to me it looks as if you do:)

OpenStudy (zzr0ck3r):

there is no asymptots for cos(x) and sin(x) there is however for sin(x)/x for the same reason as 1/x

OpenStudy (zzr0ck3r):

graph or think about 1/x note x can never equal 0 but it can get as close as it does, should read graph or think about 1/x note x can never equal 0 but it can get as close as it wants.

OpenStudy (anonymous):

i think you all don't know how to do it ...

OpenStudy (zzr0ck3r):

lol ok go with that and go figure it out and get back to us, as you have just broke math if you find it:)

OpenStudy (zzr0ck3r):

we have given you the answers you just need to think about it.

OpenStudy (zzr0ck3r):

think about 1/.0000000000000000000001 what does that equal?

OpenStudy (zzr0ck3r):

now add 50 0's

OpenStudy (zzr0ck3r):

this is the cause of such asymptotes

OpenStudy (anonymous):

your answers are not making sense with the question which i had written .... you all are telling me the basic definition of asymptote

OpenStudy (zzr0ck3r):

because you ask a question that does not make sense. It is sort of like me asking you to tell me when the function f(x) = 3 is equal to 0. it just never is....

OpenStudy (zzr0ck3r):

go graph them...wolfram alpha will do it for free:)

OpenStudy (anonymous):

@zzr0ck3r my question doesn't make sense ... what ?? go...... first study asymptotes of trigonometric functions .. then give your suggestions...

OpenStudy (anonymous):

i think this website is just for students of lower grades not for ..

OpenStudy (anonymous):

@Hero ... .where are u? how u got 99 smartscore?

OpenStudy (zzr0ck3r):

lol, listen like 5 really smart people have told you what is going on. you refuse to listen. Not all trig functions have asymptotes such as sin(x) (does not) while sin(1/x) does.

OpenStudy (zzr0ck3r):

im pretty sure your just trolling at this point...so good luck with all that. when you want to learn about math let us know. gl

OpenStudy (anonymous):

how u got sin x does not have asymptote?

OpenStudy (zzr0ck3r):

yes and notice you can draw that whole graph without lifting your pen, thus no asymtote. go graph sin(x) then graph sin(1/x) note the difference of the two and how sin(1/x) shoots of to infinity and sin(x) does not.

hero (hero):

Instead of complaining, try to understand what an asymptote is first.

OpenStudy (zzr0ck3r):

amen, again good luck. And again, those dudes ^^^^ with the 99's really seem to know what they are talking about:)

OpenStudy (anonymous):

i am disappointed...

OpenStudy (zzr0ck3r):

This has to be a joke.

OpenStudy (anonymous):

its true...

OpenStudy (anonymous):

Tell me pls what is the problem - with your "humor" i mean